OpenStudy (anonymous):

square root of 75 over 2 plus 3 square root of 5 over 4?

OpenStudy (anonymous):

\[75=25\times 3\] \[\sqrt{75}=\sqrt{25}\times \sqrt{3}=5\sqrt{3}\] is a start

OpenStudy (anonymous):

so you have \[\frac{5\sqrt{3}}{2}+\frac{3\sqrt{5}}{4}\]

OpenStudy (anonymous):

if you want to add you need the same denominator so write \[\frac{10\sqrt{3}}{4}+\frac{3\sqrt{5}}{4}\] \[=\frac{10\sqrt{3}+3\sqrt{5}}{4}\]

OpenStudy (anonymous):

no like terms in the numerator so nothing combines